Chlorine has 2 naturally occurring isotopes, 35Cl (isotopic mass 34.9689 amu) and 37Cl (isotopic mass 36.9659 amu). If chlorine has an atomic mass of 35.4527 amu, what is the percent abundance of each isotope?


1
Expert's answer
2021-05-05T01:42:44-0400

According to the equation:

(M1)(x) + (M2)(1-x) = M(E)

where M1 - the mass of one isotope, x is the relative abundance, M2 is the mass of the second isotope, M(E) is the atomic mass of the element. From here:

M1x +M2 - M2x = M(E)

x(M1 - M2) = M(E) - M2

x = (M(E) - M2) / (M1 - M2) = (35.4527 amu - 36.9659 amu) / (34.9689 amu - 36.9659 amu) = 0.7577 = 75.77 %

The abundance of 35Cl = 76 %, the abundance of 37Cl = 100 % - 75.77 % = 24.23 %.


Answer: 75.77 %; 24.23 %.